### v2

# Feynman integrals at large loop order and the $\log$-$\Gamma$ distribution

Abstract
We find empirically that the value of Feynman integrals follows a $\log$-$\Gamma$ distribution at large loop order. This result opens up a new avenue towards the large-order behavior in perturbative quantum field theory. Our study of the primitive contribution to the scalar $\phi^{4}$ beta function in four dimensions up to 17 loops provides accompanying evidence. Guided by instanton considerations, we discuss the extrapolation of this contribution to all loop orders.
1 Introduction
Feynman integrals are the building blocks of perturbative quantum, statistical, and classical field theory expansions. Each Feynman integral contributes to a specific perturbative order. In most cases, this order equals the graph’s loop number. A perturbative expansion is a formal power series $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, with the $L$-th coefficient given by a sum of Feynman integrals:
$A_{L}=\sum_{L(G)=L}\frac{I_{G}}{|\operatorname{Aut}(G)|},$ (1)
where $\hbar$ is the perturbative expansion parameter, we sum over all Feynman graphs $G$ of specific shape and loop order $L(G)$, $I_{G}$ is the Feynman integral corresponding to the graph, and $|\operatorname{Aut}(G)|$ denotes the symmetry factor of the graph (i.e. the order of its automorphism group). The precise shape of the graphs and the associated integral depend on the specific underlying theory.
The type of perturbative expansion (1) allows the prediction of a large variety of physical phenomena. For instance, the Feynman amplitude in various quantum field theories is typically expanded via (1) (see, e.g., [1]). The critical exponents of various interesting universality classes can be computed from similar expressions (e.g., for the $3$-dimensional Ising model and $3$-dimensional percolation theory [2]). Even general relativity corrections to the Newton potential can be computed using Feynman integral sums as the one above (see, e.g., [3]).
Much progress has been made in computing the coefficients $A_{L}$ in recent decades (e.g., [4, 5, 6, 7]). Improvements in understanding the underlying mathematical structures of amplitudes, Feynman integrals, and their singularities enabled these leaps [8, 9, 10, 11].
Regarding the ubiquitous nature of Feynman integrals and sums, we ask: What is the distribution of the values contributing to (1)? Are all $I_{G}$ of the same magnitude, or do particular graphs contribute more? We thus study the distribution of the value of Feynman integrals.
Ultimately, this question is motivated by the perturbation theory at large order program [12, 13] and its recent incarnation, the resurgence program (see, e.g., [14]). The progress from both these programs suggests that the large-$L$ behavior of the coefficients $A_{L}$ encodes much (and perhaps all) nonperturbative information of the function $\mathcal{A}(\hbar)$. Further, an ongoing research program aims to replace sums as (1) with integrals over a single, continuous object [15, 16, 17, 18]. Our results provide concrete information on such an object’s expected asymptotic $L\rightarrow\infty$ shape.
This article will focus on a specific model: $\phi^{4}$ quantum field theory in four-dimensional spacetime. Within this model, we focus on a particular observable: the primitive contribution to the beta function. The $\phi^{4}$ beta function is known exactly up to loop order $7$ in the minimal subtraction (MS) scheme [19, 20, 21, 4]. The primitive contribution to this beta function is obtained by summing over all period Feynman integrals, given by the $1/\varepsilon$ residues of specific Feynman integrals (see §2 for a precise definition of period Feynman integrals). It is conjectured that asymptotically, at large loop order, the primitive contribution gives the dominant contribution to the beta function of $\phi^{4}$ theory in the MS scheme [22] (see also [23, 24, 25]).
In §2, we use an empirical, numerical approach to study the terms in the sum (1) when $L$ is large. At sufficiently large loop order, exact computation methods for Feynman integrals will inevitably fail. Here, we use the tropical sampling approach introduced by the first author in [26] to evaluate many Feynman integrals with up to 17 loops (see also [27]). The tropical sampling method draws from previous ideas of sector decomposition [28, 29] and the Hepp bound [30]. Balduf and Balduf–Shaban recently performed similar large-scale computations of Feynman integrals using the tropical approach [31, 32]. In [31], the primitive contribution to the $\phi^{4}$ theory beta function was estimated up to 18 loops. We confirm these computations up to 17 loops.
The focal point of this article is the distribution of the values in the sum (1). Our main result is such a (conjectured) limiting distribution of period Feynman integrals in $\phi^{4}$ theory for large $L$. The simplicity of our result suggests similar structures within other observables in more elaborate quantum field theories. The following histogram illustrates this limiting distribution:
Figure 1: Distribution of $\phi^{4}$ period Feynman integrals at 17 loops.
The histogram shows our measured distribution of $\phi^{4}$ period Feynman integrals at 17 loops. We obtained this distribution by evaluating $44027$ Feynman integrals up to $10^{-3}$ relative accuracy. Histograms similar to Figure 1 already appeared in [33] (see also [34]), which studied the distribution of Feynman integrals with a related underlying motivation. Here, thanks to more advanced tools, we have the advantage of being able to probe a much higher order in perturbation theory. Below, Figure 2 shows similar histograms for lower loop orders and illustrates the rapid convergence to the limiting distribution.
The red curve depicts the density
$\mu(\mathcal{P})=\frac{\lambda^{\alpha}}{\Gamma(\alpha)}\left(\log\frac{\mathcal{P}}{\mathcal{P}_{0}}\right)^{\alpha-1}\left(\frac{\mathcal{P}}{\mathcal{P}_{0}}\right)^{-\lambda}\text{d}(\log\mathcal{P}),$ (2)
where $\mu(\mathcal{P})$ is the density of period Feynman integrals with value $\mathcal{P}$, and the parameters $\alpha=6.41(5)$, $\lambda=3.87(3)$, and ${{\mathcal{P}_{0}}=5.14(7)\cdot 10^{4}}$ are fitted and specific to $L=17$ (see §3.1 for details on the fitting method). The density (2) is called a $\log$-$\Gamma$ distribution because, as a function of $\log\mathcal{P}$, it is the shifted integrand of Euler’s representation for the $\Gamma$ function. This motivates
Conjecture 1.
The distribution of period Feynman integrals in $\phi^{4}$ theory at $L$ loops weakly converges to a $\log$-$\Gamma$ distribution when $L\rightarrow\infty$.
We expect this conjecture to hold for broader classes of Feynman integrals than just period Feynman integrals in $\phi^{4}$ theory. For instance, the extrapolation of the findings of [33] by analogy to our results suggests that similar limiting distributions can be observed in QED.
In general, the coefficients $A_{L}$ are expected to grow factorially [35] and, hence, the perturbative series, such as $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, to be divergent if $\hbar\neq 0$. To still utilize the information of the coefficients $A_{L}$ and to rigorously compute a quantity such as $\mathcal{A}(\hbar)$ for nonzero $\hbar$, resummation techniques are needed. As input data, these techniques need information about the $L\rightarrow\infty$ asymptotic behavior of $A_{L}$. For this reason, this asymptotic behavior is of high conceptual and practical value for the perturbative quantum field theory framework [36, 37, 38, 39, 40]. One concrete application is the prediction of critical exponents of various interesting universality classes via the Wilson–Fischer approach [2] (see also [41] for a recent discussion of the relevance of large-order contributions in this domain). In §3, we tabulate our 17 loop order results for the value of the primitive contribution to the $\phi^{4}$ beta function. We then extrapolate these values to the limit $L\rightarrow\infty$ and discuss the relation to classic conjectures on the asymptotic growth rate of the $\phi^{4}$ beta function in the MS scheme (e.g. to [42, 22, 24, 25]). We conclude in §4.
2 Methodology
2.1 The primitive contribution to the $\phi^{4}$ beta function
In four dimensions, the superficial degree of divergence of a scalar Feynman graph is given by $\omega(G)=|E_{G}|-2L(G)$, where $|E_{G}|$ is the number of edges of the graph and $L(G)$ its loop number [43]. A graph $G$ is primitive divergent in $\phi^{4}$ theory if it is 1PI, has precisely four external legs, and $\omega(\gamma)>0$ for each proper subgraph $\gamma\subsetneq G$, while $\omega(G)=0$. The last conditions ensure the graph has an overall logarithmic divergence and no subdivergences in four-dimensional spacetime. As usual, we will consider the legs of the graphs fixed or equivalently distinguishable.
The momentum representation of an $L$-loop Feynman integral in $D$ dimensions reads
$I_{G}=\frac{1}{\pi^{LD/2}}\int\frac{\text{d}^{D}k_{1}\cdots\text{d}^{D}k_{L}}{\prod_{e\in E_{G}}Q_{e}},$ (3)
where $Q_{e}=q_{e}^{2}-m_{e}^{2}+i0$ is the Feynman propagator associated with an edge $e$, and we integrate over $L$ copies of Minkowski space. We will assume that the external kinematics are sufficiently generic, so there are no IR divergences.
If $G$ has $L$ loops, $G$ is primitive divergent in a QFT that is renormalizable in four dimensions and $D=4-2\varepsilon$, then $I_{G}$, as a function of $\varepsilon$, has a simple pole at $\varepsilon=0$. The value of the associated residue is independent of the external kinematics associated to the graph, i.e.,
$I_{G}=\frac{\mathcal{P}(G)}{\varepsilon L}+\mathcal{O}(\varepsilon^{0})\text{ as }\varepsilon\rightarrow 0,$ (4)
and $\mathcal{P}(G)$ is a number independent of the external kinematics. We call the number $\mathcal{P}(G)$ the period Feynman integral associated with graph $G$ (see, e.g., [44] for more details on periods).
Our main example is the primitive contribution to the $\phi^{4}$ beta function (see [23, Appendix B], whose notation we follow, for details). This contribution is renormalization group independent and therefore appears in every renormalization scheme. We may express this contribution in terms of period Feynman integrals,111 The convential shift by one in $\beta^{\mathrm{prim}}_{L+1}$ is due to the fact that $L$-loop vertex diagrams contribute to the $g^{L+1}$ coefficient of the beta function in the coupling $g$.
$\beta^{\mathrm{prim}}_{L+1}=2\sum_{\begin{subarray}{c}G\\ \phi^{4}\text{ primitive}\\ L(G)=L\end{subarray}}\frac{\mathcal{P}(G)}{|\operatorname{Aut}G|},$ (5)
where we sum over all primitive divergent $\phi^{4}$ graphs at $L$ loops.
We aim to get a clearer picture of the behavior of this sum and its terms when $L$ is large. For $L=17$, the sum (5) has $\approx 7\cdot 10^{12}$ terms, making it impractical to evaluate all of them individually. For this reason, we use a sampling approach to study (5) and its terms.
2.2 A probabilistic approach to the sum over graphs
Instead of summing over the potentially large number of Feynman graphs at fixed loop order $L$, we will sample such graphs $G$ with probability
$p(G)=\frac{1}{Z_{L}}\frac{1}{|\operatorname{Aut}G|},$ (6)
where the normalization factor $Z_{L}$ is given by
$Z_{L}=\sum_{\begin{subarray}{c}G\\ \phi^{4}\text{ primitive}\\ L(G)=L\end{subarray}}\frac{1}{|\operatorname{Aut}G|}.$
The following algorithm generates samples of primitive divergent graphs with probability (6):
Algorithm 1 Generate random primitive $L$-loop $\phi^{4}$ graph
1:Start with $L+1$ isolated vertices that each have four distinguishable legs.
2:Randomly select two of the legs and connect them, replacing two legs with one new edge.
3:Repeat the last step until only four legs are left.
4:If the resulting graph is primitive divergent, then return the graph. If not, go back to step 1.
As the legs are distinguishable (which can be realized on the computer by numbering them, for example), the graph returned by the algorithm will have four distinguishable legs. It follows from a simple combinatorial argument and the orbit stabilizer theorem that the algorithm above randomly produces graphs $G$ with probability $p(G)$ from (6). Combining (5) and (6), gives
$\beta^{\mathrm{prim}}_{L+1}=2\cdot Z_{L}\cdot\sum_{\begin{subarray}{c}G\\ \phi^{4}\text{ primitive}\\ L(G)=L\end{subarray}}p(G)\cdot\mathcal{P}(G)=2\cdot Z_{L}\cdot\langle\mathcal{P}(G)\rangle_{L},$ (7)
where we recover the expectation value $\langle\mathcal{P}(G)\rangle_{L}$ of the random variable $\mathcal{P}(G)$ under the probabilistic process over the set of $L$-loop primitive divergent graphs described by Algorithm 1.
We, therefore, can study the sum (5) by sampling graphs using this algorithm. We do so as follows: We first generate a sample of a primitive divergent graph using Algorithm 1. Then, we evaluate the associated period Feynman integral using the tropical sampling algorithm from [26] and the implementation [45] (see the second author’s Master’s thesis [46] for a gentle introduction to the tropical integration method). We configured this tropical sampling algorithm to compute the value of each sampled Feynman graph to about $10^{-3}$ relative accuracy. The resulting number provides one data point for an evaluated period Feynman integral. We repeat these steps a large number of times. For example, Figure 1 summarizes all 44027 data points we obtained at 17 loops by running Algorithm 1 and the tropical sampling algorithm the same number of times.
3 Results
3.1 Histograms of Feynman integrals at large loop order
We used the methods described in the last section to generate representative samples of primitive divergent $\phi^{4}$ graphs and evaluate their period Feynman integrals at loop orders $8$ to $17$. The number of graphs we sampled at each loop order is listed in Table 1. We ran the computation in bunches at low priority on the ETH Euler computing cluster. Due to a maintenance event, our computation was interrupted, and some data points were lost. Hence, the number of samples differs slightly at each loop order. There is no correlation between the probability of a data point being lost and its value. At 17 loops, we took fewer samples because we only had limited access to the required large-memory nodes. All our data points of randomly sampled Feynman graphs are available as machine-readable tables in the ancillary material to this article’s arXiv version.
Figure 1 and Figure 2 depict the sampled period Feynman integrals as histograms. We evaluated each Feynman graph to $10^{-3}$ relative accuracy using the tropical sampling approach. As this uncertainty is small compared to the statistical uncertainty that stems from the variance of the different Feynman graphs, we can neglect this uncertainty. We confirmed this explicitly by performing our analysis with the uncertainty included and obtaining identical results.
Figure 2: Distribution of $\phi^{4}$ period Feynman integrals up to 16 loops. The $x$-axes show the value of $\log(\mathcal{P})$, and the $y$-axes show the density.
Our data suggests that the distribution of period Feynman integrals is modelled well by the distribution (2) for $L\rightarrow\infty$. At each loop order, we fitted the parameters $\alpha,\lambda,$ and $\mathcal{P}_{0}$ by maximizing the logarithmic likelihood function
$\log\mathcal{L}=N\left(\alpha\log\lambda-\log\Gamma(\alpha)\right)+(\alpha-1)\sum_{i}\log\log\frac{\mathcal{P}_{i}}{\mathcal{P}_{0}}-\lambda\sum_{i}\log\frac{\mathcal{P}_{i}}{\mathcal{P}_{0}},$
where we sum over all period samples $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}$ at a specific loop order. This likelihood function is readily derived by multiplying $N$ copies of the function that multiplies the measure in (2) and taking the logarithm. As the number of samples $N$ is large, we can estimate the uncertainties of these parameters by approximating the prior distribution using a Gaussian. The resulting parameters with uncertainties are listed in Table 1. The uncertainties of the fit parameters were extremely large for $L\in\{8,9\}$. So, we discarded these fits. The fitted distributions are depicted as red lines in Figure 1 and Figure 2.
We checked Conjecture 1 quantitatively using Pearson’s $\chi^{2}$ test: Let $O_{i}$ be the number of evaluated period Feynman integrals that fall into the $i$-th percentile of the distribution (2) with the fitted maximum likelihood parameters at the respective loop order. The expectation value of the random variable $O_{i}$ is $N/100$. So, under the hypothesis that our data follows (2), the quantity $\chi^{2}=\frac{100}{N}\sum_{i}(O_{i}-\frac{N}{100})^{2}$ is expected to follow a $\chi^{2}$-distribution with mean $100-3=97$, as three parameters are fitted. Table 1 shows that the ratio $\chi^{2}/97$ approaches $1$ with increasing loop order, consistent with Conjecture 1. Figure 2 illustrates how the distribution is approached with increasing loop number, providing further evidence for Conjecture 1.
$L$ $\alpha$ $\lambda$ $\mathcal{P}_{0}$ $\frac{\chi^{2}}{97}$ $\langle\mathcal{P}(G)\rangle_{L}$ $\beta_{L+1}^{\rm prim}$    $N$
$8$ $7.449(7)\cdot 10^{2}$ $6.064(5)\cdot 10^{6}$ $99900$
$9$ $1.644(2)\cdot 10^{3}$ $1.046(1)\cdot 10^{8}$ $99800$
$10$ $19.72(84)$ $10.6(2)$ $4.97(19)\cdot 10^{2}$ $273.7$ $3.504(5)\cdot 10^{3}$ $1.890(3)\cdot 10^{9}$ $99900$
$11$ $10.69(27)$ $6.9(1)$ $1.36(3)\cdot 10^{3}$ $67.7$ $7.237(12)\cdot 10^{3}$ $3.558(6)\cdot 10^{10}$ $99700$
$12$ $8.36(17)$ $5.53(6)$ $2.78(4)\cdot 10^{3}$ $19.3$ $1.460(3)\cdot 10^{4}$ $6.998(14)\cdot 10^{11}$ $100000$
$13$ $7.38(13)$ $4.85(5)$ $5.28(6)\cdot 10^{3}$ $7.3$ $2.869(6)\cdot 10^{4}$ $1.429(3)\cdot 10^{13}$ $98712$
$14$ $6.88(11)$ $4.42(4)$ $9.58(10)\cdot 10^{3}$ $4.6$ $5.531(14)\cdot 10^{4}$ $3.036(7)\cdot 10^{14}$ $100000$
$15$ $6.83(10)$ $4.24(4)$ $1.67(2)\cdot 10^{4}$ $2.5$ $1.039(3)\cdot 10^{5}$ $6.654(18)\cdot 10^{15}$ $100000$
$16$ $6.688(96)$ $4.08(4)$ $2.92(3)\cdot 10^{4}$ $2.2$ $1.910(6)\cdot 10^{5}$ $1.504(4)\cdot 10^{17}$ $99650$
$17$ $6.41(12)$ $3.87(5)$ $5.14(7)\cdot 10^{4}$ $1.2$ $3.489(17)\cdot 10^{5}$ $3.556(17)\cdot 10^{18}$ $44027$
Table 1: Fit parameters $\alpha,\lambda,$ and $\mathcal{P}_{0}$ for the period Feynman integral distribution in $\phi^{4}$ theory, the normalized $\chi^{2}$ value resulting from Pearson’s $\chi^{2}$ test, the estimated average value of the period Feynman integral, the estimated value of the primitive beta function coefficient, and the number $N$ of Feynman integrals evaluated at each loop order.
3.2 Estimating moments and the primitive $\phi^{4}$ beta function
The results of [31] suggest that the moments $\langle\mathcal{P}(G)^{k}\rangle_{L}$ diverge for $L\rightarrow\infty$ if $k\geq 2$ (see Table 8 and the discussion before Eq. (4.8) loc. cit.). Hence, for sufficiently large loop order, the central limit theorem, which requires a finite second moment, might not be applicable to estimate the value of $\langle{\mathcal{P}}(G)\rangle_{L}$. So, we can expect a naive sampling approach to fail eventually, as it will become unfeasible to estimate $\langle{\mathcal{P}}(G)\rangle_{L}$ in Eq. (5) via averaging over a relatively small amount of samples. However, under the assumption of the validity of Conjecture 1, we can compute $\langle{\mathcal{P}}(G)\rangle_{L}$ for large $L$ directly using Eq. (2), provided that the parameters $\alpha,\lambda,$ and $\mathcal{P}_{0}$ are known:
$\langle{\mathcal{P}}(G)^{k}\rangle_{L}=\frac{\lambda^{\alpha}}{\Gamma(\alpha)}\int_{\mathcal{P}_{0}}^{\infty}\mathcal{P}^{k}\left(\log\frac{\mathcal{P}}{\mathcal{P}_{0}}\right)^{\alpha-1}\left(\frac{\mathcal{P}}{\mathcal{P}_{0}}\right)^{-\lambda}\text{d}(\log\mathcal{P})=\mathcal{P}_{0}^{k}\left(\frac{\lambda}{\lambda-k}\right)^{\alpha}\text{ for large }L.$ (8)
The integral only converges if $\lambda>k$, so outside this range, we find $\langle{\mathcal{P}}(G)^{k}\rangle_{L}=\infty$.
In our data (see Table 1), we observed that $\lambda>3$ for $L\leq 17$. Hence, up to this loop order, we may assume that $\langle\mathcal{P}(G)^{2}\rangle_{L}$ is finite and that we can safely estimate $\langle\mathcal{P}(G)\rangle_{L}$ using the central limit theorem by computing the average of all computed values of $\mathcal{P}(G)$. The central limit theorem has the advantage over Eq. (8) of providing accurate results if the data does not exactly follow distribution (2). The uncertainty of this estimate is computed as usual by dividing the sample variance by the number of samples and taking the square root. The results of the estimation are listed in Table 1. We also estimated $\langle{\mathcal{P}}(G)\rangle_{L}$ using Eq. (8). The discrepancy between both ways to estimate this expectation value decreases with the loop order $L$. At 17 loops, both methods give confidence intervals that overlap almost perfectly.
Via (7), we can translate our estimates for $\langle\mathcal{P}(G)\rangle_{L}$ into the primitive contribution to the $\phi^{4}$ beta function. The necessary values of $Z_{L}$ were calculated using renormalized 0-dimensional QFT technology [47, §6.3] or [48] (see also [49] for a deeper analysis of similar normalization factors from 0-dimensional quantum field theory). Table 1 also includes the resulting estimates for $\beta_{L+1}^{\rm prim}$. These agree with the data for $L\leq 11$ in [23, Table XIII] and confirm estimates in [31, Table 14] up to 17 of 18 loop orders.
3.3 Extrapolating $\beta_{L}^{\rm prim}$ to all loop orders using instanton input
The coefficients $\beta_{L}^{\rm prim}$ of the primitive contribution to the $\phi^{4}$ beta function are expected to grow factorially. In fact, assuming the conjecture that $\beta^{\mathrm{prim}}_{L}\sim\beta^{\mathrm{MS}}_{L}$ for large $L$, instanton computations in scalar quantum field theories [50, 22, 24] (see also §IV.B of [23] for details) give a precise prediction for the asymptotic behavior:
$\beta^{\mathrm{prim}}_{L}=C\cdot L^{{7}/{2}}\cdot L!\cdot\left(1+\mathcal{O}\left(\frac{1}{L}\right)\right),$ (9)
where $C$ equals
$\frac{144\cdot e^{-\frac{15}{4}-3\,\gamma_{E}}}{\pi^{3/2}A^{6}}\approx 0.024199,$
with $\gamma_{E}$ Euler’s constant, and $A$ the Glaisher–Kinkelin constant (Eq. (21) of [23] with $n=1$).
Knowing the precise asymptotic behavior of quantities such as $\beta^{\mathrm{MS}}_{L}$ for $L\rightarrow\infty$ is essential to fruitfully apply resummation techniques while turning divergent perturbative expansions into trustworthy predictions [38, 39, 23, 40]. Our results give new hard data on this behavior.
Besides the primitive contribution to this asymptotic behavior, renormalons are expected to give a factorially growing contribution in the $L\rightarrow\infty$ limit. Computing the renormalon contributions to the $\phi^{4}$ beta function in the MS scheme and analyzing their asymptotic growth rate seems achievable using, e.g., technology from [51, 52, 53, 54, 55]. Comparing these contributions with the primitive contribution would shed new light on the question of whether instanton or renormalon contributions dominate at large loop order [56, 57, 25, 41].
Unfortunately, our data for $\beta^{\mathrm{prim}}_{L}$ does not provide compelling evidence for the asymptotic behavior (9). We find, via goodness-of-fit estimates, that our data for $\beta^{\mathrm{prim}}_{L}$ is compatible with the following Ansatz, which Eq. (9) directly inspires:
$\beta^{\mathrm{prim}}_{L}=L^{{7}/{2}}\cdot L!\cdot\left(c_{0}+\frac{c_{1}}{L}+\frac{c_{2}}{L^{2}}+\cdots\right)\text{ as }L\rightarrow\infty\,,$ (10)
where $c_{0},c_{1},\ldots$ are free parameters. Fitting this Ansatz to our data for $\beta^{\mathrm{prim}}_{L}$ results in decent fits. Fixing the fit parameter $c_{0}$ to $C$, as defined above, reduces the fit quality dramatically.
By varying the fit range and the cut-off point of the expansion in powers of $1/L$ in (10), we obtained various estimates for the value of $c_{0}$. With reasonable choices for both the fit range and the cut-off point, $c_{0}$ consistently falls into the range of $0.055\pm 0.015$. Even though the predicted value of $C\approx 0.024$ lies outside this range, the large margin of error does not allow for a clear verdict on the validity of (9). Moreover, the results of [49] show that asymptotic estimates such as ours, which rely on low-order computations, suffer from various biases. The authors highlight a particular case relevant to $\phi^{4}$ theory, where the true asymptotic behavior only becomes apparent if the first 25 perturbative coefficients are known.
4 Conclusion
Empirically, we observed that the value of Feynman integrals converges (weakly) to a specific distribution once the loop order gets large. We provided evidence for this by studying the primitive contribution to the $\phi^{4}$ beta function in four-dimensional spacetime. The specific limiting distribution we find is the $\log$-$\Gamma$ distribution — a distribution well-known in statistics and probability theory. We expect limiting distributions of the same or similar shapes to appear in Feynman perturbative expansions of observables in other quantum field theories.
The limiting distribution has three remaining parameters, $\alpha,\lambda,$ and $\mathcal{P}_{0}$, that we fix by fitting at each loop order. It would be highly beneficial to find explicit limiting laws for these three parameters, i.e. to find the asymptotic behavior of these numbers for $L\rightarrow\infty$.
We gathered large amounts of data on the value of Feynman integrals to come to our conclusions. This data is available with the arXiv version of this article.
We used our data to compute the primitive contribution to the $\phi^{4}$ beta function up to 17 loops confirming previous results by Balduf [31]. We discussed the extrapolation of this contribution to all loop orders. It is conjectured that the primitive beta function equals the $\phi^{4}$ beta function in the minimal subtraction scheme when $L\rightarrow\infty$, so our extrapolation provides a (conjectured) estimate of also this beta function at infinite loop order. This second beta function has numerous phenomenological applications via the Wilson–Fischer approach to critical phenomena. We postpone the discussions of the phenomenological implications of our findings (e.g., on the critical exponents of the $D=3$ Ising model) to future work.
Unfortunately, reaching a verdict of full or only partial agreement between our data and predictions from instanton computations is still impossible. Our limited data, the resulting poor fit quality, and the large number of sources for numerical perturbations do not allow a clear conclusion. More data at an even higher loop order seems necessary to complete the picture. The key limiting factor of our computations was the memory requirements of the tropical sampling implementation [45]. Harnessing more properties of the Hepp bound from [30] and the (tropical) geometry of Feynman integrals might reduce these requirements and make higher loop orders accessible (see [26, §8.1 and §8.3]).
Acknowledgments
We thank Paul Balduf, Erik Panzer, and Oliver Schnetz for their valuable comments and suggestions for an early version of this paper, Paul Balduf and Erik Panzer for helpful discussions, and Babis Anastasiou for his support. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. MB was supported by Dr. Max Rössler, the Walter Haefner Foundation and the ETH Zürich Foundation. This work was supported by the Swiss National Science Foundation through its project funding scheme (grant number 10001706). We used the ETH Euler cluster for our calculations.